# Conditional entropy and information of quantum processes

### v1

@@ VII.1 Quantum hypothesis testing for channel discrimination
The generalized conditional entropy $\mathbf{S}[A|B]_{\mathcal{N}}$ of a bipartite channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ has an operational meaning in terms of the channel discrimination task [57, 58, 34, 59] based on the associated generalized divergence $\mathbf{D}[\cdot\|\cdot]$ between the given channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ (null hypothesis) and some channel ~$\widetilde{\mathcal{R}}_{A^{\prime}\to A}\otimes\mathcal{M}_{B^{\prime}\to B}$ (alternate hypothesis) (cf. [34, Appendix D]), where ~$\widetilde{\mathcal{R}}_{A^{\prime}\to A}:=\frac{1}{|A|}\mathcal{R}_{A^{\prime}\to A}$ is the completely mixing or depolarizing channel,
$-\mathbf{S}[A|B]_{\mathcal{N}}$ $=\inf_{\mathcal{M}\in{\mathrm{Q}}(B^{\prime},B)}\mathbf{D}[\mathcal{N}_{A^{\prime}B^{\prime}\to AB}\|\mathcal{R}_{A^{\prime}\to A}\otimes\mathcal{M}_{B^{\prime}\to B}],$ (104)
@@
Operational meanings and comparison of von Neumann conditional entropy, min conditional entropy, and geometric hypothesis testing relative entropy in terms of asymmetric hypothesis testing for channel discrimination can be done by invoking  [34, Theorem 49, Appendix D] and inequality (43).
To elucidate, we consider a game with two different kind of adversarial challenges $\texttt{Quest}_{x}[\mathcal{N}]$, where $\in\{a,b\}$. The game consists of a challenger, Charlie, and an adversary, Ash. Ash provides a black box $\mathcal{B}$, which is a bipartite quantum channel $\mathcal{B}^{y}$ ($y\in\{0,x\}$), to Charlie and the goal of Charlie is to discriminate if the given box indeed is the desired bipartite channel. In each challenge $x$, black box has only two possibilities $\{H_{0},H_{1}^{x}\}$.
@@ 2. $H_{1}^{x}$
      if $x=a$, $\mathcal{B}_{A^{\prime}B^{\prime}\to AB}^{a}=\widetilde{\mathcal{R}}_{A^{\prime}\to A}\otimes\mathcal{M}_{B^{\prime}\to B}$, where a channel $\mathcal{M}_{B^{\prime}\to B}$ is chosen in an idealized way that would give best mimicry of $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ in a given setting where the completely mixing channel ~$\widetilde{\mathcal{R}}_{A^{\prime}\to A}$ is fixed.
@@
obtained from the $\varepsilon$-hypothesis testing relative entropy $D^{\varepsilon}_{h}(\rho\|\sigma):=-\log\beta_{\varepsilon}(\rho\|\sigma)$ and $\beta_{\varepsilon}:=\inf_{0\leq\lambda\leq\mathbbm{1}}\{\operatorname{tr}[\lambda\sigma]:\operatorname{tr}[\lambda\rho]\geq 1-\varepsilon\}$.
@@
    $R^{\parallel}_{b}[\mathcal{N}]=\inf_{\mathcal{M}\in{\mathrm{Q}}(B^{\prime},B)}D[\mathcal{N}_{A^{\prime}B^{\prime}\to AB}\|\widetilde{R}_{A\to A}\circ\mathcal{N}_{A^{\prime}B^{\prime}\to AB}]=-S^{\not\to}[A|B]_{\mathcal{N}}+\log|A.$ (111)
If a quantum channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}\in\mathrm{NS}_{A^{\prime}\not\to B}$, then $R^{\parallel}_{a}[\mathcal{N}]=R^{\parallel}_{b}[\mathcal{N}]$ always holds. There exists bipartite channels $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}\notin\mathrm{NS}_{A^{\prime}\not\to B}$ for which $R^{\parallel}_{a}[\mathcal{N}]>R^{\parallel}_{b}[\mathcal{N}]$.
@@ VII.2 Conditional quantum channel merging
The von Neumman entropy $S[A]_{\mathcal{M}}$ of a quantum channel $\mathcal{M}_{A^{\prime}\to A}$ is equal to the quantum channel merging capacity [19], which is a variant of quantum state merging protocol. We note that the entropy of a channel is related to its private randomness capacity $P_{\mathrm{r}andom}(\mathcal{M})$ (up to a constant factor of log of the output dimension) [21, Theorem 6]; to be precise, $P_{\mathrm{r}andom}(\mathcal{M})+S[A]_{\mathcal{M}}=\log|A|$. The goal of quantum channel merging is to transfer the state of output $A$ to Eve possessing isomteric extension $E$ of the channel $\mathcal{M}$ [19], where $\mathcal{V}^{\mathcal{M}}_{A^{\prime}\to AE}$ is an isometric channel extension of $\mathcal{M}$. We consider a variant of the quantum channel merging protocol, which we call conditional quantum channel merging protocol.
@@
where $\Phi^{l}_{A^{\prime\prime}_{1}E^{\prime\prime}_{1}}$ and $\Phi^{k}_{A^{\prime\prime}_{0}E^{\prime\prime}_{0}}$ are maximally entangled states of Schmidt rank $l$ and $k$, respectively, and system pairs $A^{\prime\prime}_{0},A^{\prime}_{1}$ and $E^{\prime\prime}_{0},E^{\prime\prime}_{1}$ are auxiliary systems of Ash and Eve, respectively.
The entanglement gain of the protocol is $\log M:=\log l-\log k$ and $R:={\log M}/{n}$ is called the entanglement (gain) rate of the protocol. $R$ is called achievable rate if there exists a protocol $(n,2^{n(R-\delta)},\varepsilon)~{}~{}\forall~{}~{}\epsilon\in(0,1],\delta>0$ for sufficiently large $n$. The conditional quantum channel merging capacity $C_{\mathrm{g}ain}(\mathcal{N}_{A^{\prime}B^{\prime}\rightarrow AB})$ is defined as
@@
The conditional quantum channel merging capacity of an arbitrary bipartite channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ is upper bounded by the
@@
where $S^{\not\to}[A^{n}|B^{n}]_{\mathcal{N}^{\otimes n}}$ is the von Neumann conditional NS-entropy of a bipartite channel $\mathcal{N}$.
We omit the detailed proof here as the proof argument for the above theorem is similar (almost same arguments) to the (converse) proof of [19, Proposition 23], which is closely based on [14, Section IV.B.]. To arrive at the inequality (115), we made the observation that
@@
For any tensor-product channel $\mathcal{N}^{1}_{A^{\prime}\to A}\otimes\mathcal{N}^{2}_{B^{\prime}\to B}$, $C_{\mathrm{g}ain}[A|B]_{\mathcal{N}^{1}\otimes\mathcal{N}^{2}}=S^{\not\to}[A|B]_{\mathcal{N}^{1}\otimes\mathcal{N}^{2}}=S[A|B]_{\mathcal{N}^{1}\otimes\mathcal{N}^{2}}=S[A]_{\mathcal{N}^{1}}$, due to Theorem 2 and [19, Theorem 10].
VIII Strong subadditivity for channels
@@
Tripartite quantum states $\rho_{ABC}$ for which $I(A;C|B)_{\rho}=0$ are called quantum Markov states [63, 64].
@@
where $\Phi_{R_{A}AR_{B}BR_{C}C}=\mathcal{N}(\Phi_{R_{A}A^{\prime}}\otimes\Phi_{R_{B}B^{\prime}}\otimes\Phi_{R_{C}C^{\prime}})$.
@@
Given that $\frac{1}{2}\norm{\mathcal{N}-\mathcal{T}}_{\diamond}=\varepsilon$, we have $\frac{1}{2}\norm{\Phi^{\mathcal{N}}-\Phi^{\mathcal{T}}}_{1}\leq\varepsilon$. We then arrive at the bound (127) by directly invoking Theorem 6 and the continuity of quantum conditional mutual information (see [65, Corollary 2]): For two states $\rho_{ABC},\sigma_{ABC}$ such that $\frac{1}{2}\norm{\rho-\sigma}_{1}=\varepsilon\in[0,1]$ then
@@ VIII.1 Approximate Quantum Markov channel
For an arbitrary quantum state $\rho_{ABC}$, there exists a universal recovery map $\mathcal{R}_{B\to BC}$ such that [66, 64] (also see [67, Theorem 2])
@@
A tele-covariant channel $\mathcal{T}_{A^{\prime}B^{\prime}C^{\prime}\to ABC}$ is simulable via teleportation with the resource state as its Choi state $\Phi^{\mathcal{T}}_{R_{A}AR_{B}BR_{C}C}$ [9, Theorem 5]. That is, the action of the tele-covariant channel $\mathcal{T}$ can be obtained by local (quantum) operations and classical communication (LOCC) among Ash (holding $R_{A},A^{\prime},A$), Bob (holding $R_{B},B^{\prime},B$), and Charlie (holding $R_{C},C^{\prime},C$), if Ash, Bob, and Charlie share the Choi state $\Phi^{\mathcal{N}}$ among themselves,
@@
We succinctly present the findings from the above observation as a lemma below (see also [23, Theorem 2 and Proposition 9] for a broader statement).

### v2

@@ VII.1 Quantum hypothesis testing for channel discrimination
The generalized conditional entropy $\mathbf{S}[A|B]_{\mathcal{N}}$ of a bipartite channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ has an operational meaning in terms of the channel discrimination task [60, 61, 37, 62] based on the associated generalized divergence $\mathbf{D}[\cdot\|\cdot]$ between the given channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ (null hypothesis) and some channel ~$\widetilde{\mathcal{R}}_{A^{\prime}\to A}\otimes\mathcal{M}_{B^{\prime}\to B}$ (alternate hypothesis) (cf. [37, Appendix D]), where ~$\widetilde{\mathcal{R}}_{A^{\prime}\to A}:=\frac{1}{|A|}\mathcal{R}_{A^{\prime}\to A}$ is the completely mixing or depolarizing channel,
$-\mathbf{S}[A|B]_{\mathcal{N}}$ $=\inf_{\mathcal{M}\in{\mathrm{Q}}(B^{\prime},B)}\mathbf{D}[\mathcal{N}_{A^{\prime}B^{\prime}\to AB}\|\mathcal{R}_{A^{\prime}\to A}\otimes\mathcal{M}_{B^{\prime}\to B}]$
@@
Hence, we have $\mathbf{S}[A|B]_{\mathcal{N}}=\log|A|-\inf_{\mathcal{M}\in{\mathrm{Q}}(B^{\prime},B)}\mathbf{D}[\mathcal{N}_{A^{\prime}B^{\prime}\to AB}\|\widetilde{\mathcal{R}}_{A^{\prime}\to A}\otimes\mathcal{M}_{B^{\prime}\to B}]$ if $\mathbf{D}(\rho\|c\sigma)=-\log c+\mathbf{D}(\rho\|\sigma)$ for all scalars $c\in\mathbb{C}$.
VII.1.1 Parallel uses of channels for the channel discrimination task
To elucidate, we consider a game with two different kind of adversarial challenges $\texttt{Quest}_{x}[\mathcal{N}]$, where $x\in\{a,b\}$. The game consists of a challenger, Charlie, and an adversary, Ash. Ash provides a black box $\mathcal{B}$, which is a bipartite quantum channel $\mathcal{B}^{y}$ ($y\in\{0,x\}$), to Charlie and the goal of Charlie is to discriminate if the given box indeed is the desired bipartite channel. In each challenge $x$, black box has only two possibilities $\{H_{0},H_{1}^{x}\}$.
@@ 2. $H_{1}^{x}$
      if $x=a$, $\mathcal{B}_{A^{\prime}B^{\prime}\to AB}^{a}=\widetilde{\mathcal{R}}_{A^{\prime}\to A}\otimes\mathcal{M}_{B^{\prime}\to B}$, where a channel $\mathcal{M}_{B^{\prime}\to B}$ is chosen in the best possible way to help adversary mimic $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ given that the completely mixing channel ~$\widetilde{\mathcal{R}}_{A^{\prime}\to A}$ is fixed.
@@
obtained from the $\varepsilon$-hypothesis testing relative entropy $D^{\varepsilon}_{h}(\rho\|\sigma):=-\log\beta_{\varepsilon}(\rho\|\sigma)$ and $\beta_{\varepsilon}(\rho\|\sigma):=\inf_{0\leq\lambda\leq\mathbbm{1}}\{\operatorname{tr}[\lambda\sigma]:\operatorname{tr}[\lambda\rho]\geq 1-\varepsilon\}$.
@@
    $R^{\parallel}_{b}[\mathcal{N}]=D[\mathcal{N}_{A^{\prime}B^{\prime}\to AB}\|\widetilde{R}_{A\to A}\circ\mathcal{N}_{A^{\prime}B^{\prime}\to AB}]=-S^{\not\to}[A|B]_{\mathcal{N}}+\log|A|.$ (119)
If a quantum channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}\in\mathrm{NS}_{A^{\prime}\not\to B}$, then $R^{\parallel}_{a}[\mathcal{N}]=R^{\parallel}_{b}[\mathcal{N}]$ always holds. In contrast, for the bipartite channels $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}\notin\mathrm{NS}_{A^{\prime}\not\to B}$ we have $R^{\parallel}_{a}[\mathcal{N}]>R^{\parallel}_{b}[\mathcal{N}]$.
@@
VII.1.2 Adaptive strategy for channel discrimination
Asymmetric hypothesis testing for channel discrimination using adaptive strategy.— Let $\{\Lambda,\mathcal{A}\}$ denote an arbitrary adaptive protocol for discriminating quantum channels $\mathcal{N}^{1},\mathcal{N}^{2}\in{\mathrm{Q}}(A^{\prime},A)$, where $\Lambda$ denotes the measurement operator corresponding to the null hypothesis at the end of the protocol and $\mathcal{A}$ denotes the set of (adaptive) channels applied between the uses of $\mathcal{N}^{1},\mathcal{N}^{2}$; see [55, Section 3] for the definitions of the protocol and also for the (slightly modified) notations used here. Let $\alpha_{n}\left(\{\Lambda,\mathcal{A}\}\right)$ and $\beta_{n}\left(\{\Lambda,\mathcal{A}\}\right)$ be the corresponding type-I and type-II error probabilities, respectively, where $n\in\mathbb{N}$ denotes the number of channel uses for $\mathcal{N}^{1}$ or $\mathcal{N}^{2}$ in the protocol. The asymmetric distiguishability, in Stein’s setting, is defined as [55]
$\zeta_{n}\left(\varepsilon,\mathcal{N}^{1},\mathcal{N}^{2}\right):=\sup_{\Lambda,\mathcal{A}}\bigg{\{}-\frac{1}{n}\log\beta_{n}\left(\{\Lambda,\mathcal{A}\}\right)\big{|}\alpha_{n}\left(\{\Lambda,\mathcal{A}\}\right)\leq\varepsilon\bigg{\}},$ (121)
and its asymptotic quantities are defined as
$\mkern 2.0mu\overline{\mkern-2.0mu\zeta\mkern-2.0mu}\mkern 2.0mu\left(\varepsilon,\mathcal{N}^{1},\mathcal{N}^{2}\right)=\underset{n\rightarrow\infty}{\lim\sup}~{}\zeta_{n}\left(\varepsilon,\mathcal{N}^{1},\mathcal{N}^{2}\right),\ \text{and}\ \underaccent{\bar}{\zeta}\left(\varepsilon,\mathcal{N}^{1},\mathcal{N}^{2}\right)=\underset{n\rightarrow\infty}{\lim\inf}~{}\zeta_{n}\left(\varepsilon,\mathcal{N}^{1},\mathcal{N}^{2}\right).$ (122)
Let $\varepsilon\in(0,1)$ and $\alpha\in(1,2]$, then we have the following strong converse bound on $\mkern 2.0mu\overline{\mkern-2.0mu\zeta\mkern-2.0mu}\mkern 2.0mu\left(\varepsilon,\mathcal{N}^{1},\mathcal{N}^{2}\right)$  [37, Theorem 49, Appendix D] (also see  [55, Corollary 18]),
$D\left(\mathcal{N}^{1}\|\mathcal{N}^{2}\right)\leq\underaccent{\bar}{\zeta}\left(\varepsilon,\mathcal{N}^{1},\mathcal{N}^{2}\right)\leq\mkern 2.0mu\overline{\mkern-2.0mu\zeta\mkern-2.0mu}\mkern 2.0mu\left(\varepsilon,\mathcal{N}^{1},\mathcal{N}^{2}\right)\leq\widehat{D}_{\alpha}\left(\mathcal{N}^{1}\|\mathcal{N}^{2}\right)\leq D_{\mathrm{m}ax}\left(\mathcal{N}^{1}\|\mathcal{N}^{2}\right).$ (123)
The input state for an adaptive discrimination protocol can be arbitrary. The above bound works for any pair of quantum channels $\mathcal{N}^{1},\mathcal{N}^{2}\in{\mathrm{Q}}(A^{\prime},A)$.
Applying inequalities (123) and (45), for any bipartite quantum channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$,
$-S[A|B]_{\mathcal{N}}+\log|A|$ $\leq\inf_{\mathcal{M}\in{\mathrm{Q}}(B^{\prime},B)}\underaccent{\bar}{\zeta}\left(\varepsilon,\mathcal{N}_{A^{\prime}B^{\prime}\to AB},\widetilde{\mathcal{R}}_{A^{\prime}\to A}\otimes\mathcal{M}_{B^{\prime}\to B}\right)$ (124)
$\leq\inf_{\mathcal{M}\in{\mathrm{Q}}(B^{\prime},B)}\mkern 2.0mu\overline{\mkern-2.0mu\zeta\mkern-2.0mu}\mkern 2.0mu\left(\varepsilon,\mathcal{N}_{A^{\prime}B^{\prime}\to AB},\widetilde{\mathcal{R}}_{A^{\prime}\to A}\otimes\mathcal{M}_{B^{\prime}\to B}\right)$ (125)
$\leq-\widehat{S}_{\alpha}[A|B]_{\mathcal{N}}+\log|A|\quad\forall\alpha\in(1,2]$ (126)
$\leq-S_{\mathrm{m}in}[A|B]_{\mathcal{N}}+\log|A|.$ (127)
@@ VII.2 Conditional quantum channel merging
The von Neumman entropy $S[\mathcal{M}]$ of a point-to-point quantum channel $\mathcal{M}_{A^{\prime}\to A}$ is equal to the quantum channel merging capacity [20]. The goal of quantum channel merging is to transfer the state of output $A$ to Eve possessing isomteric extension $E$ of the channel $\mathcal{M}$ [20], where $\mathcal{V}^{\mathcal{M}}_{A^{\prime}\to AE}$ is an isometric channel extension of $\mathcal{M}$. It is a variant of a quantum state merging protocol. The entropy of a channel is equal to its completely bounded entropy $S_{\mathrm{C}B,min}$ [66, Eq. (1.1)]. We note that the entropy of a channel is related to its private randomness capacity $P_{\mathrm{r}andom}(\mathcal{M})$ (up to a constant factor of log of the output dimension) [22, Theorem 6]. The private randomness capacity is the optimal rate, in an usual asymptotic $i.i.d.$ uses of channels, at which the receiver (Bob) accessing output $A$ of the channel $\mathcal{M}_{A^{\prime}\to A}$ can extract private randomness, against eavesdropper accessing isomteric extension $E$ of the channel ${\mathcal{N}}_{A^{\prime}\to A}$, when the sender (Ash) sends states through the channel $\mathcal{N}$ (see [22] for proper definition). In particular, we have the following complimentary relation between $P_{\mathrm{r}andom}(\mathcal{M})$ and $S[\mathcal{M}]$,
$P_{\mathrm{r}andom}(\mathcal{M})+S[\mathcal{M}]=\log|A|.$ (128)
The sum of the private randomness capacity $P_{\mathrm{r}andom}(\mathcal{M})$ and the entropy of a channel $S[A]_{\mathcal{M}}$ always equals to the logarithm of the output dimension of the channel. The private randomness capacity is maximum, $2\log|A|$, if and only if the channel is unitary and minimum, 00, if and only if the channel is completely mixing (depolarizing).
We consider a variant of the quantum channel merging protocol, which we call conditional quantum channel merging protocol.
@@
where $\Phi^{l}_{A^{\prime\prime}_{1}E^{\prime\prime}_{1}}$ and $\Phi^{k}_{A^{\prime\prime}_{0}E^{\prime\prime}_{0}}$ are maximally entangled states of Schmidt rank $l$ and $k$, respectively, and system pairs $A^{\prime\prime}_{0},A^{\prime}_{1}$ and $E^{\prime\prime}_{0},E^{\prime\prime}_{1}$ are auxiliary systems of Ash and Eve, respectively. The entanglement gain of the protocol is $\log M:=\log l-\log k$ and $R:=\frac{1}{n}{\log M}$ is called the entanglement (gain) rate of the protocol. $R$ is called an achievable rate if there exists a protocol $(n,2^{n(R-\delta)},\varepsilon)~{}~{}\forall~{}~{}\epsilon\in(0,1],\delta>0$ for sufficiently large $n$. The conditional quantum channel merging capacity $C_{\mathrm{g}ain}(\mathcal{N}_{A^{\prime}B^{\prime}\rightarrow AB})$ is defined as
@@
The conditional quantum channel merging capacity of an arbitrary bipartite channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ is upper bounded (weak converse) by the
@@
where $\frac{1}{n}S^{\not\to}[A^{n}|B^{n}]_{\mathcal{N}^{\otimes n}}$ is the regularized von Neumann conditional NS-entropy, i.e., $S^{\not\to}[A^{n}|B^{n}]_{\mathcal{N}^{\otimes n}}$ is the von Neumann conditional NS-entropy of the $n$ parallel uses of a bipartite channel $\mathcal{N}$.
We omit the detailed proof here as the proof argument for the above theorem is similar (almost same arguments) to the (converse) proof of [20, Proposition 23], which is closely based on [15, Section IV.B.]. To arrive at the inequality (131), we made the observation that
@@
For any tensor-product channel $\mathcal{N}^{1}_{A^{\prime}\to A}\otimes\mathcal{N}^{2}_{B^{\prime}\to B}$, $C_{\mathrm{g}ain}[A|B]_{\mathcal{N}^{1}\otimes\mathcal{N}^{2}}=S^{\not\to}[A|B]_{\mathcal{N}^{1}\otimes\mathcal{N}^{2}}=S[A|B]_{\mathcal{N}^{1}\otimes\mathcal{N}^{2}}=S[A]_{\mathcal{N}^{1}}$, due to Theorem 2 and [20, Theorem 10].
VIII Strong subadditivity of the entropy of channels
@@
The structure of states that satisfy the strong subadditivity condition with equality were studied in [67]. Tripartite quantum states $\rho_{ABC}$ for which $I(A;C|B)_{\rho}=0$ are called quantum Markov states [67, 68]. In this case there exist an operation from subsystem $B$ to $BC$ which can reconstruct $\rho_{ABC}$ from $\rho_{AB}$.
@@
where $\Phi^{\mathcal{N}}_{R_{A}AR_{B}BR_{C}C}=\mathcal{N}(\Phi_{R_{A}A^{\prime}}\otimes\Phi_{R_{B}B^{\prime}}\otimes\Phi_{R_{C}C^{\prime}})$.
@@
Given that $\frac{1}{2}\norm{\mathcal{N}-\mathcal{T}}_{\diamond}=\varepsilon$, we have $\frac{1}{2}\norm{\Phi^{\mathcal{N}}-\Phi^{\mathcal{T}}}_{1}\leq\varepsilon$. We then arrive at the bound (143) by directly invoking Theorem 6 and the continuity of quantum conditional mutual information (see [69, Corollary 2]): For two states $\rho_{ABC},\sigma_{ABC}$ such that $\frac{1}{2}\norm{\rho-\sigma}_{1}=\varepsilon\in[0,1]$ then
@@ VIII.1 Approximate Quantum Markov channel
For an arbitrary quantum state $\rho_{ABC}$, there exists a universal recovery map $\mathcal{R}_{B\to BC}$ such that [70, 68] (also see [71, 72] for relevant discussions)
@@
A tele-covariant channel $\mathcal{T}_{A^{\prime}B^{\prime}C^{\prime}\to ABC}$ is simulable via teleportation with the resource state as its Choi state $\Phi^{\mathcal{T}}_{R_{A}AR_{B}BR_{C}C}$ [10, Theorem 5]. That is, the action of the tele-covariant channel $\mathcal{T}$ can be obtained by local (quantum) operations and classical communication (LOCC) among Ash (holding $R_{A},A^{\prime},A$), Bob (holding $R_{B},B^{\prime},B$), and Charlie (holding $R_{C},C^{\prime},C$), if Ash, Bob, and Charlie share the Choi state $\Phi^{\mathcal{N}}$ among themselves,
@@
We succinctly present the findings from the above observation as a lemma below (see also [24, Theorem 2 and Proposition 9] for a broader statement).